OpenStudy (anonymous):

Please help! Simplify the product!! (7x3y2z) (7xy4z2)

OpenStudy (saifoo.khan):

(7x^3 y^2 z) (7 x y^4 z^2) Open the like terms up! 7*7 = 49 x^3 * x = x^4 y^2* y^4 = y^6 z * z^2 = z^3

OpenStudy (saifoo.khan):

The property we use, \[\Large a^b\times a^c=a^{b+c}\]

OpenStudy (australopithecus):

add the exponents on like variables remember that 32xyb is just one number we just dont know what x, y or b is equal to also remember that x^(1) is the same as x 2^(2) = 2 2^(2) * 2^(2) = 2^(4) = 16

OpenStudy (saifoo.khan):

Therefore, \[\Large 49x^4 y^6 z^3\]

OpenStudy (anonymous):

thank you! I really appreciate it! :)